OpenStudy (anonymous):

What percent of 67 is 33? If necessary, round to the nearest tenth of a percent. A. 203.0% B. 0.5% C. 49.3% D. 33.0%

OpenStudy (anonymous):

Let the percentage by x: x % of 67 is 33, So, \[\frac{x}{100} \times 67 = 33 \implies x = \frac{33 \times 100}{67}\] Solve this to find x..

OpenStudy (anonymous):

its c right?

OpenStudy (anonymous):

Absolutely..

OpenStudy (anonymous):

thank you very much i knew the formula was something close to that just didnt know how to write it

OpenStudy (amistre64):

i learned these using a picture